OpenStudy (anonymous):

Consider the equation an^2=182 where a is any number between 2 and 5 and n is a positive integer. What are the possible values of n???

OpenStudy (anonymous):

I only found 9

OpenStudy (diyadiya):

an^2 = 182 Divide both sides by a n^2 = 182/a n = \(\sqrt{182/a}\) cont..

OpenStudy (hoblos):

an^2=182 n=√(182/a) so you must find the values of a such that n stays between 2 and 5 for a=8 n=4.77 accepted for a=45 n=2.01 accepted for (a) less than 8 n>5 for (a) greater than 45 n<2

OpenStudy (hoblos):

and for each value of a between 8 and 45 you have a different value of n between 2 and 5

OpenStudy (anonymous):

great

OpenStudy (hoblos):

do you need all the possible values or just the number of possible values??

OpenStudy (diyadiya):

a is any number between 2 & 5 182/5=36.4 182/2=91 Since √182/a is an integer 182/a is a perfect square (between 36.4 & 91) there are 3 perfect square 49 ,64 ,81 between 36.4 and 91 therefore the values of n are \[\sqrt{49}=7 \ \sqrt{64}=8 \ \sqrt{81}=9\]

OpenStudy (anonymous):

@hoblos all the possible values.

OpenStudy (callisto):

Diya, you still have the -ve values to include~ like -7x-7 =49

OpenStudy (hoblos):

n is a (positive) integer @Callisto

OpenStudy (callisto):

Sorry... I was blind :(

OpenStudy (diyadiya):

No sorry =D

OpenStudy (anonymous):

so what (n)*49=182????

OpenStudy (diyadiya):

its a*49=182 a is any number b/w 2&5

OpenStudy (anonymous):

yes sorry, what a??

OpenStudy (diyadiya):

You can check the answer a*49=182 a=182/49= 3.7 (approx.)

OpenStudy (diyadiya):

3.7 is a number between 2 & 5

OpenStudy (anonymous):

Fantastic.

OpenStudy (diyadiya):

=)